Attachment 'runinfo_from_elog_TBMar2018.txt'

Download

   1 12.03.2018
   2 new modules are installed in the testbeam 24. 
   3 5 HBU6 modules are connected to wing-LDA. 
   4  
   5 layer	module	LDA port	chip ID
   6 1	22	14	33-48
   7 2	26	15	209-224
   8 3	23	16	241-0
   9 4	24	17	1-16
  10 5	25	26	17-32
  11 
  12 Noise runs
  13 7006: no validation, no beam (2000 cycles - enough) 
  14 7007: validation on, no beam (6000 cycles - just enough) 
  15  
  16 DAQ statistics:  
  17 7006: 28 ROC/s, 45% duty cycle 
  18 7007: 42 ROC/s, 72% duty cycle  
  19 
  20 Runs with absorber
  21 7008-7011: Validation off 
  22 7012: Validation off, event building off (however still with EB in the data collector, which matches 
  23  everything with BIF - more work to be done to disable this step) 
  24 7013: Validation on, overnight run with temperature compensation on. 20 minutes per run 
  25 
  26 Beam parameters
  27 3 Gev, target 2 (selected and displayed). 
  28 Collimator in the beam area: 8x8mm 
  29 Collimators:  
  30     +5.3 
  31 +5.4    +5.3 
  32     +5.4 
  33 Moving stage counters reset to the center of the layers 
  34 Beam rate: 3.5 kHz 
  35 
  36 13.03.2018
  37 Overnight runs
  38 7013-7044 running happily 
  39 
  40 start of MIP scan
  41 15k triggers.  
  42 runs 7045-7072 (stopped manually) 
  43 
  44 Timing issue
  45 1 layers showed different timing, than others.  
  46 Power cycled, absorber in. 
  47 run 7073 - shooting to the center with absorber, validation on 
  48 run 7074 - shooting to the center with absorber, validation off 
  49  
  50 Reprogrammed DIF on 1st layer, after that: 
  51 7075 (val off, center, absorber in) 
  52 7076 (val on, center, absorber in) 
  53 
  54 Offsets
  55 New offsets: AHCAL: 2101, BIF:67493 
  56 Previous runs were taken with a wrong offset(2123, 68110) 
  57 
  58 MIP scan
  59 started from run 7078 
  60 
  61 run stop 
  62 at 7125 
  63 position: x-353 y307 
  64 
  65 restart run
  66 start run 7126 
  67 position: x-353 y307 
  68 
  69 EUDAQ crash
  70 BIF producer crashed, blocked the whole DAQ. 
  71 Restarted from y-23, x-23 
  72 
  73 ISSUE
  74 3rd layers, 13th asic (counting from 1), ch 31 has a weird electrical crosstalk. Whenever there 
  75  is a hit around, this channel has often a trigger 
  76 
  77 switch to absorber runs
  78 Interrupted MIP scan at y=-173, x=-353, run7513 
  79 
  80 overnight run with absorber
  81 run 7514-7537 
  82 
  83 14.03.2018
  84 
  85 Absorber removed. Resumed position at Y=-173, X=-353 
  86 Run 7538- :  
  87 
  88 Short LED scan 7541XX during the beam maintenance  
  89 Continuing scan from 7541 
  90 
  91 Started MIP scan  
  92 run number 7702 
  93 module number 5,14,15,32,40 
  94  
  95 layer	module	LDA port	chip ID
  96 5(1)	5	14	177-192
  97 14(2)	14	15	145-160
  98 15(3)	15	16	161-176
  99 32(4)	32	17	49-64
 100 40(5)	40	26	65-80
 101 
 102 stop run at 7966 
 103 position x337, y7 
 104 
 105 Al Absorber in beam line
 106 start from 7967 at x0 , y0 
 107 
 108 
 109 15.03.2018
 110 
 111 overnight run successfully finished with run 7990 
 112 run without the validation: 7991 
 113 
 114 Removed absorber, continue with MIP scan 
 115 position y=7, x=337, run 7992 
 116 Beam energy 3GeV 
 117 Intensity ~4kHz 
 118 
 119 BIF crashed at y-53_x-323 position. After run 8062. 
 120 EUDAQ restarted by Jiri. BIF problems are gone. 
 121 Data taking started from the same position, run number 8064. 
 122 Temperatures are below +50C 
 123 
 124 MIP scanning is over, last run number: 8306 
 125 hort LED SCAN
 126 run numbers 8316XX
 127 
 128 layer	module	LDA port	chip ID
 129 2(1)	2	14	177-192
 130 8(2)	8	15	161-176
 131 30(3)	30	16	81-96
 132 36(4)	36	17	97-112
 133 37(5)	37	26	113-128
 134 
 135 8318: noise run, no beam, validation off (HV adjustments too high) 
 136 8319: noise run, no beam, validation off (HV adjustments OK) 
 137 8320: noise run, no beam, validation off, ambient light on (to check light contamination) (see attached pdf file for hitmaps) 
 138  
 139 Start MIP scan
 140 RUN: 8329, absorber removed, first temp comp every 10min 
 141  
 142 Stop Muon Scan, Start Absorber Runs
 143 Stopped Muon Run at Run Nr: 8497 at x: 337 y: 127 
 144 Start Al. Absorber Run at Run Nr. 8498 at x: 0 y:0  
 145 
 146 16.03.2018
 147 	
 148 EUDAQ crash - out of hard drive space. 
 149 Run 8517 - last run (where the hdd space ran out) 
 150 
 151 Restart MIP scan: 
 152 Run 8519, x=337, y=127 
 153 
 154 MIP scan finished at run 8930. 
 155 we start short LED run. but module30 have a problem on slow control loading. we then switched off 
 156  HV/LV and wait 5 minutes for cooling down the module. 
 157  
 158 
 159  new modules are installed 
 160 module 18 and 28 have 680pF capacitor instead of 1nF capacitor. 
 161  
 162 layer	module	LDA port	chip ID
 163 18(1)	18	14	129-144
 164 28(2)	28	15	145-160
 165 31(3)	31	16	161-176
 166 38(4)	38	17	177-192
 167 39(5)	39	26	193-208
 168  
 169 noise run
 170 run 8936 without validation 
 171 run 8937 with validation 
 172 
 173 start MIP scan from run 8941 
 174 temperature compensation every 10 minutes 
 175 
 176 stop run
 177 		
 178 
 179 stop run at 9016 
 180 position X247 y247 
 181 
 182 
 183 17.03.2018
 184 
 185 start shower run
 186 start shower run from 9017. with absorber, validation on, temperature compensation every 30 minutes 
 187 
 188 absorber run finished at 9:00, run 9043 
 189 9044: resumed MIP scan 
 190 
 191 EUDAQ crash (Data collector dc1 during restart - TCP error). Resumed in run 9173 
 192 
 193 stop at y=-293, x=-83, run9490 
 194 
 195 
 196 18.03.2018
 197 Scan with absorber finished at run 9534 
 198 
 199 MIP scan
 200 MIP scan started from run 9535 
 201 
 202 MIP scan finished
 203 stop run 9593 
 204  
 205 short LED run 
 206 
 207 start shower run from 9594 with 5 GeV electron. 
 208 
 209 19.03.2018
 210 
 211 Out of HDD space!
 212 Run crashed on Sunday ~ 17:00 
 213 No new modules instructions: continue with 5GeV e-
 214 
 215 Runs 9614-9622: Moved to x=90, y=-90 to get large energy deposition in single ASIC 
 216 
 217 Run with absorber, hiting an asic (x=90, y=-90) 
 218 9623- 
 219 
 220 Gain Threshold Scan
 221 		
 222 
 223 Replaced 3 last modules: 
 224 layer	module	LDA port	chip ID
 225 18(1)	18	14	129-144
 226 28(2)	28	15	145-160
 227 12(3)	12	16	241-0
 228 11(4)	11	17	225-240
 229 10(5)	10	26	33-48
 230  
 231 Start Runs for Gain Threshold Scans 
 232 RUN 9637 
 233 DAC2 Gain Sel.: 500 
 234 First HBU X180 Y-180 
 235 
 236 BIF Error
 237 		
 238 
 239 During Run 9637 
 240 Repeat Run at the same position X180 Y-180 
 241 -> RUN 9639 
 242 
 243 Error Data Receiver
 244 		
 245 
 246 DataReceiver:  Unable to stop the data receiving/forwarding... 
 247 After RUN 9639 was finished, Data ok? 
 248 BIF ok this time. 
 249 
 250 Restart LabVIEW 
 251 Continue Gain Threshold Scan 
 252 For HBU 2  Position: X -180 Y -180 
 253 RUN: 9641 
 254 immediately crashed due to Error  
 255 restart Eudaq and continue with RUN 9642 
 256 looks fine now 
 257 
 258 Restart Gain Threshold Scan
 259 Run: 9644 
 260 Restart gain threshold scan at central position X0 Y0 for 300000 events 
 261 
 262 Continue over night, RUN 9645 onwards 
 263 
 264 20.03.2018
 265 
 266 restart gain threshold scan
 267 run 9670: position Y180 X-180, VALIDATION not working! bad data 
 268 run 9671: position Y180 X-180 
 269 run 9672: position Y180 X180 
 270 run 9673: position Y0,X0 
 271 
 272 Threshold 450
 273 SC programming problems with layer 10 again. Cooling down 
 274 9674: position x180, y-180 
 275 9675: position x-180, y-180 
 276 9676: position  x-180, y+180 
 277 9677: position x+180, y+180 
 278 9678-9679: position 0,0 
 279 
 280 thr 400
 281 cooldown for reconfiguration was necessary again. 
 282 9680: x+180 y-180 
 283 9681: x-180 y-180 
 284 9682: x-180 y+180 
 285 
 286 extra 8cm absorber added in fron of the 27 cm absorber. Absorber moved closer to the detector 
 287 9685: gain thr 400, position x0, y0  
 288 
 289 
 290 21.03.2018
 291 
 292 beam delivery stopped in run 9709 
 293 
 294 Auto trigger test
 295 9712 - test with 15 dummy triggers.  
 296 LED=6750, Enable section: Trig_ext synchr off, TCALIB1(5) on, TCALIB_direct(9) on 
 297 
 298 5GeV beam, Threshold 350
 299 Runs: 
 300 9714: X+180, Y-180 
 301 9715: X-180, Y-180 
 302 9716: X-180, Y+180 
 303 9717: X+180, Y+180 
 304 9718: X0, Y0 
 305 9719?????: X+15, Y+15 
 306 
 307 	
 308 Plan to test today: 
 309 RUN: 9721, 9722 *test without validation (x+15,y+15) 
 310 RUN:9720 *move the beam to the center of 1 tile to get higher deposition (x+15,y+15) 
 311     move the beam to the center of 1 tile to get higher deposition (x-15,y-15)
 312 RUN 9723 *test with 2 dummy triggers (x+15,y+15) 
 313     test with LED AT dummy trigger
 314     Increase the trigger threshold
 315 RUN 9724 *test with 2 dummy triggers (x-15,y-15) 
 316 RUN 9725- *move the beam to the center of 1 tile to get higher deposition (x-15,y-15) 
 317  
 318 
 319 22.03.2018
 320 
 321 9725-9743: pos x-15, y-15. overnight 
 322  
 323 9744-9745: trying LED AT with 1 dummy trigger (trig_ext off), position x=0,y=0 
 324 9746-0747: 1AT dummy, Aim at channel 35 (x-15,y+15) 
 325 9748: 1ATdummy, val off, position [0,0] 
 326 9749: 2ATdummy, val off, position [0,0] 
 327 9755: 2AT, val off, Aim at channel 35 (x-15,y+15) 
 328 
 329 reprogrammed for trigger threshold 280, Gain sel thr 350. position x-15,y+15 
 330 Run: 9757-9761 
 331 
 332 1 AT trig, tthr 280, gthr 350, val off: 
 333 9763: x15,y-15 (channels 5) 
 334 9764 x15,y15 (channel 30) 
 335 9765: x-15,y15 (channel 35) 
 336 9766: x0y0 (stopped) 
 337 
 338 switching to ag400, threshold 280. SC reload failed in modules 12 and 10. Cooldown 
 339 LED 5950, position x-15, y15 (chan 35), 2 dummy using tcalib_hold 
 340 Runs 9770-9780  
 341 
 342 23.03.2018
 343 
 344 switching to ag400, threshold 280. SC reload failed in modules 12 and 10. Cooldown 
 345 LED 5950, 2 dummy using tcalib_hold 
 346 Run 9781: x-15, y-15 
 347 -9785 (0,0) 
 348 
 349 With validation: 
 350 run 9786: x-15, y-15 
 351 
 352 
 353 
 354 24.03.2018

Attached Files

To refer to attachments on a page, use attachment:filename, as shown below in the list of files. Do NOT use the URL of the [get] link, since this is subject to change and can break easily.
  • [get | view] (2018-04-13 08:17:32, 87.8 KB) [[attachment:MapChipChan2IJK_Mar2018_batch1_2.txt]]
  • [get | view] (2018-04-13 08:17:56, 44.3 KB) [[attachment:MapChipChan2IJK_Mar2018_batch3.txt]]
  • [get | view] (2018-04-13 08:18:18, 44.9 KB) [[attachment:MapChipChan2IJK_Mar2018_batch4.txt]]
  • [get | view] (2018-04-13 08:18:43, 44.3 KB) [[attachment:MapChipChan2IJK_Mar2018_batch5.txt]]
  • [get | view] (2018-04-13 08:18:55, 44.9 KB) [[attachment:MapChipChan2IJK_Mar2018_batch6.txt]]
  • [get | view] (2018-03-23 12:50:21, 8.5 KB) [[attachment:runinfo_from_elog_TBMar2018.txt]]
 All files | Selected Files: delete move to page copy to page

You are not allowed to attach a file to this page.